# Learning to Swim: Reinforcement Learning for 6-DOF Control of Thruster-driven Autonomous Underwater Vehicles

### v1

@@ Abstract
Controlling AUVs can be challenging because of the effect of complex non-linear hydrodynamic forces acting on the robot, which, unlike ground robots, are significant in water and cannot be ignored. The problem is especially challenging for small AUVs for which the dynamics can change significantly with payload changes and deployments under different water conditions. The common approach to AUV control is a combination of passive stabilization with added buoyancy on top and weights on the bottom, and a PID controller tuned for simple and smooth motion primitives. However, the approach comes at the cost of sluggish controls and often the need to re-tune controllers with configuration changes. We propose a fast (trainable in minutes), reinforcement learning based approach for full 6 degree of freedom (DOF) control of an AUV, enabled by a new, highly parallelized simulator for underwater vehicle dynamics. We demonstrate that the proposed simulator models approximate hydrodynamic forces with enough accuracy that a zero-shot transfer of the learned policy to a real robot produces performance comparable to a hand-tuned PID controller. Furthermore, we show that domain randomization on the simulator produces policies that are robust to small variations in vehicle’s physical parameters.
@@ I INTRODUCTION
Various control methodologies currently exist for AUVs including PID, sliding mode [3], and model predictive control [4]. However, when considering how AUVs are operating in environments that are varying across space and time, it becomes necessary for control methods to be capable of adapting to unpredictable changes in hydrodynamics. Furthermore, when deploying AUVs in the field, operators may wish to attach various payloads to the vehicle, shifting its physical parameters. On smaller AUVs, this can be especially problematic as payloads or small changes in environmental parameters (such as water density from salinity) can have large effects on the hydrodynamics (such as the center of buoyancy and mass) that are difficult to model and measure. When using classical methods such as PID, this can result in a time-consuming tuning process where the controller must be modified to support the new AUV configuration. Additionally, it is common to increase passive stability of the vehicle in these contexts by adding float to the top and weights on the bottom of the vehicle, but this trades of controllability and agility. Thus, a more robust and adaptive method of controlling AUVs is needed to support more streamlined or sophisticated field deployments.
In this work we contribute: (1) a highly-parallelized, GPU-based AUV simulator designed for learning robotic controls using reinforcement learning methods; (2) we also show how our simulator can be used to evaluate the robustness of RL-based controls to varying physical parameters; (3) finally, we demonstrate how controls learned in our simulator can be effectively deployed on a real vehicle with performance comparable to PID control.
[width=]figures/qual-overview.png
@@ II-A Reinforcement Learning for Control
[width=]figures/fig-system-overview.png
On-policy reinforcement methods can often be trained efficiently in simulation, minimizing wear and tear on vehicles for training purposes [14]. However, because simulation environments do not perfectly model the real-world, there is often a simulation-to-real performance (Sim2Real) gap. By leveraging a strategy known as domain randomization (DR), which simulates uncertainty in dynamics and environmental parameters [15] many systems are able to immediately apply to a variety of environmental contexts robustly and without tuning [13].
@@
Hadi et al. [17] explored the use of deep RL for learning 2-DOF control (yaw, speed) in a simulator. Masmitja et al. [18] used RL to develop a tracking behavior for a surface vehicles, enabling it to follow underwater moving targets, but only focused on high-level steering commands. Lu et al. [19] used domain randomization to improve RL based control of a 4-DOF underwater robot.
@@ III-C1 Learning environment and algorithm
At a high-level, our approach is outlined in Figure 2. In this work, we explore the use of a highly parallelized simulation environment coupled with an on-policy learning algorithm that is capable of quickly utilizing the experience gained. Furthermore, as proposed in [13, 14], we utilize domain randomization, in which various aspects of the simulated environment can be sampled randomly, to train a controller that is agnostic to those environmental changes (up to a point).
@@ III-C2 Observations, actions, and rewards
where is the difference between the desired position and current position of the vehicle, in the vehicle’s local frame of reference. As in [32], we learn a controller that drives this offset to the origin, which can then be applied for either position holding or trajectory following, while avoiding the need to sample the entire unbounded state space. We clip the offset error in deployments by the amount seen during training to avoid spurious behaviors. are orientation quaternions and are the angular velocities currently measured. The policy network outputs 6 thruster commands as actions, , that are normalized between {-1, 1}, which are scaled to PWMs. During training, environments are initialized at uniformly random across possible unit rotations and random offsets from respective origins of up to 2-meters, and are also provided random goal unit rotations. Resets occur at fixed step counts, equivalent to 3 seconds of experience time. We also use the guidance term presented in [32] to speed up training.
We define a positive reward function as follows:
@@ III-C3 Domain transfer (validation in simulation)
[width=]figures/fig-neural-disturb-qual.png
[width=]figures/target_v_obs_real_disturb_exp_neural.png
[width=]figures/target_v_obs_real_disturb_exp_pid.png
@@ III-D Domain transfer (sim-to-real)
To test the transfer-ability of the trained network, we deploy our selected policy from simulation onto CUREE [5]. CUREE is an AUV equipped with 6 BlueRobotics T200 thrusters for 6-DOF movement and a Nvidia Jetson Orin NX, in a water-cooled housing, that allows us to run our neural network policy on-board in real-time ( 20Hz). On-board cameras (which are used to detect AprilTags), a Doppler Velocity Logger (DVL) and an Inertial Measurement Unit (IMU) are fused via an EKF to provide position, orientation, and velocity estimates in the body frame Figure 2. The sensors and thrusters are all controlled and monitored through a Raspberry Pi 4, and everything is implemented in ROS for coordination. Furthermore CUREE is small enough to fit in checked luggage and is actively deployed in various coral reef, marine, and freshwater environments [5]. These attributes make CUREE an ideal platform for testing our RL-based controls.
The policy takes state estimates from the EKF and a desired pose from a user or trajectory, and outputs thruster commands. We note that unlike in [32, 33], we do not model motor action delays which can lead to instabilities in deployment. For this, we apply a quaternion slerp between the desired and estimated quaternions for stable sim-to-real transfer. To test the robustness and accuracy of the controller, CUREE is commanded to hold a position above an AprilTag while it is aggressively perturbed via external forces. This setup is shown in Figure 7.
@@ IV-C Sim-to-real transfer
We test zero-shot transfer of the policy directly on real hardware in a small tank environment (which limits our abilities to test full trajectories). This version of CUREE has various sensors such as hydrophones and extra computers on a different housing, none of which were measured or integrated into the model. We are thus deploying on a vehicle with mostly unknown hydrodynamics. By recording both the commanded poses and estimated poses we are able to evaluate how closely CUREE is able to hold a position and how it recovers from aggressive disturbances as shown in Figure 8 and Figure 7. These show that CUREE, with the trained neural network policy, running at roughly 20Hz, is able to reject aggressive disturbances. This is rather surprising given the simplistic simulated modeling and domain randomization. When compared to the performance of the PID controller Figure 9, the neural controller seems more aggressive, which aligns with [13, 32].
Additionally, it is interesting to note the steady-state error that is present in the neural controller, especially in the pitch and y-directions in Figure 8. This is because the controller, as currently configured, only learns a conservative policy that works well on average, but cannot adapt to specific deployment conditions. In the future, online learning and adaptation techniques such as those presented in [33] can help alleviate these types of errors.
Further errors that we noticed in deployments that are not captured by our model are issues such as major thruster imbalances (housings that do not have streamlined geometry severely hinder thruster performance in assymmetric ways).
@@ V Conclusion
In this work, we provided a hydrodynamics implementation in Isaac Lab that enables highly parallelized simulations for underwater vehicles. We use it to train a control policy that maps high-level 6-DOF commands directly into thruster commands, in minutes, that is able to zero-shot transfer to various dynamics settings immediately. We further validate, in real-world experiments, that it can run in real-time on-board a custom AUV which enables it to recover from aggressive disturbances even when provided with full 6-DOF commands.
In the future we plan to explore using online learning and adaptation techniques such as those proposed by [34, 33, 35, 36] to develop a controller that can adapt to new physical circumstances in real-time. We hope to integrate these with our simulation in order to warmstart the model and then further fine-tune it through online learning to provide domain adaptation and resolve issues such as the steady state errors we observed in our previous experiments.
@@ ACKNOWLEDGMENTS
Thanks to Ethan Fahnestock for the initial IsaacSim-based simulator implementation and reinforcement learning pipeline. Thanks to the WHOI Summer Student Fellowship Program for financial support. Thanks also to Nathan McGuire, Peter Werner, Daniel Yang, and Tim Seyde for their help and discussions.

### v2

@@ Abstract
Controlling AUVs can be challenging because of the effect of complex non-linear hydrodynamic forces acting on the robot, which are significant in water and cannot be ignored. The problem is exacerbated for small AUVs for which the dynamics can change significantly with payload changes and deployments under different hydrodynamic conditions. The common approach to AUV control is a combination of passive stabilization with added buoyancy on top and weights on the bottom, and a PID controller tuned for simple and smooth motion primitives. However, the approach comes at the cost of sluggish controls and often the need to re-tune controllers with configuration changes. In this paper, we propose a fast (trainable in minutes), reinforcement learning-based approach for full 6 degree of freedom (DOF) control of a thruster-driven AUVs, taking 6-DOF command-conditioned inputs direct to thruster outputs. We present a new, highly parallelized simulator for underwater vehicle dynamics. We demonstrate this approach through zero-shot sim-to-real (with no tuning) transfer onto a real AUV that produces comparable results to hand-tuned PID controllers. Furthermore, we show that domain randomization on the simulator produces policies that are robust to small variations in vehicle’s physical parameters.
@@ I INTRODUCTION
Currently, there are various control methodologies for AUVs, including PID, sliding mode [3], and model predictive control [4]. However, when considering how AUVs operate in environments that vary in space and time, control methods must be able to adapt to unpredictable changes in hydrodynamics. Furthermore, when deploying AUVs in the field, operators may wish to attach various payloads to the vehicle, shifting its physical parameters. On smaller AUVs, this can be especially problematic as payloads or small changes in environmental parameters (such as water density from salinity) can have large effects on the hydrodynamics (such as the center of buoyancy and mass) that are difficult to model and measure.
Classical methods tend to require either significant and complex engineering to design robust and adaptive controllers, which are difficult to implement or transfer to new vehicles, or require time and labor-intensive manual tuning in each new configuration [5]. Additionally, it is common to increase passive stability of the vehicle in these contexts by adding float to the top and weights on the bottom of the vehicle, but this trades of controllability and agility. Thus, a simpler method to design robust and adaptive controllers for AUVs is needed to support more streamlined or sophisticated field deployments.
In this work we contribute: (1) a highly-parallelized, GPU-based AUV simulator for designing and learning controls using reinforcement learning methods 111Code available at https://github.com/warplab/isaac-auv-env, (2) to our knowledge, the first real-world demonstrations of a command-conditioned 6-DOF controller, direct to thrusters, deployed zero-shot on an AUV and (3) investigations of the sim-to-real deployment gaps on AUVs addressed by domain randomization and how it compares with hand-tuned PID controllers. We believe these strategies can reduce controller development effort and ease transferability to new systems and configurations.
[width=]figures/train-deploy-overview.png
@@ II-A Reinforcement Learning for Control
[width=]figures/fig-system-overview-white-bg.png
On-policy RL methods can often be trained efficiently in simulation, minimizing wear and tear on vehicles for training purposes [15]. However, because simulation environments do not perfectly model the real-world, there is often a simulation-to-real performance (Sim2Real) gap. By leveraging a strategy known as domain randomization (DR), which simulates uncertainty in dynamics and environmental parameters [16] many systems are able to immediately apply to a variety of environmental contexts robustly and without tuning [14].
@@
Hadi et al. [18] explored the use of deep RL for learning 2-DOF control (yaw, speed) in a simulator. Masmitja et al. [19] used RL to develop a tracking behavior for a surface vehicle, enabling it to follow underwater moving targets, but only focused on high-level steering commands. Lu et al. [20] is most similar to our approach, using RL with DR for sim-to-real transfer on an AUV, but does not have high-parallelization and limited to controls up to 4-DOF commands.
@@ III-C1 Learning environment and algorithm
At a high-level, our approach is outlined in Figure 2. In this work, we explore the use of a highly parallelized simulation environment coupled with an on-policy learning algorithm that is capable of quickly utilizing the experience gained. Furthermore, as proposed in [14, 15], we utilize domain randomization, as in [14], in which various aspects of the simulated environment can be sampled randomly, to train a controller that is agnostic to those environmental changes (up to a point).
@@ III-C2 Observations, actions, and rewards
where is the difference between the desired position and current position of the vehicle, in the vehicle’s local frame of reference. As in [33], we learn a controller that drives this offset to the origin, which can then be applied for either position holding or trajectory following, while avoiding the need to sample the entire unbounded state space. We clip the offset error in deployments by the amount seen during training to avoid spurious behaviors and mitigates impacts of saturation. are orientation quaternions and are the angular velocities currently measured. The policy network outputs 6 thruster commands as actions, , that are normalized between {-1, 1}, which are scaled to PWMs. During training, environments are initialized at uniformly random across possible unit rotations and random offsets from respective origins of up to 2-meters, and are also provided random goal unit rotations. Resets occur at fixed step counts, equivalent to 3 seconds of experience time. We also use the guidance term presented in [33] to speed up training.
We define the positive reward function at each time step , which is the objective function the reinforcement learning algorithms attempts to maximize, as follows:
@@ III-C3 Domain transfer (validation in simulation)
[width=]figures/disturbance-timeline.png
[width=]figures/disturbance_rejection_neural.png
[width=]figures/disturbance_rejection_pid.png
@@ III-D Domain transfer (sim-to-real)
To test the transfer-ability of the trained network, we deploy our selected policy from simulation onto CUREE [6]. CUREE is an AUV equipped with 6 BlueRobotics T200 thrusters for 6-DOF movement and a Nvidia Jetson Orin NX, in a water-cooled housing, that allows us to run our neural network policy on-board in real-time (20Hz). On-board cameras (which are used to detect AprilTags), a Doppler Velocity Logger (DVL) and an Inertial Measurement Unit (IMU) are fused via an EKF to provide position, orientation, and velocity estimates in the body frame Figure 2. The sensors and thrusters are all controlled and monitored through a Raspberry Pi 4, and everything is implemented in ROS for coordination. These attributes make CUREE an ideal platform for testing our RL-based controls.
The policy takes state estimates from the EKF and a desired pose from a user or trajectory, and outputs thruster commands. We note that unlike in [33, 34], we do not model motor action delays which can lead to instabilities in deployment. To account for un-modeled delays, we apply a quaternion slerp to smooth out large jumps between the desired and estimated quaternions for stable sim-to-real transfer. To test the robustness and accuracy of the controller, CUREE is commanded to hold a position above an AprilTag while it is aggressively perturbed via external forces. This setup is shown in Figure 7.
@@ IV-C Sim-to-real transfer
We test zero-shot transfer of the policy directly on real hardware in a small tank environment. We are deploying on a vehicle with relatively unknown hydrodynamics, as it contains an arbitrary sensor payload configuration and we do not measure the new center of mass or buoyancy. By recording both the commanded poses and estimated poses we are able to evaluate how closely CUREE is able to hold a position and how it recovers from aggressive disturbances as shown in Figure 8 and Figure 7. These show that CUREE, with the trained neural network policy, running at roughly 20Hz, is able to reject aggressive disturbances. This is rather surprising given the simplistic simulated modeling and domain randomization. We note around the 1.4-min mark, an error causes an unplanned disturbance, but the vehicle returns to steady state. When compared to the performance of the PID controller Figure 9, the neural controller seems more aggressive, which aligns with findings in similar studies in quadrotors [14, 33].
Additionally, it is interesting to note the steady-state error that is present in the neural controller, especially in the pitch and y-directions in Figure 8. This is because the controller only learns a conservative policy that works well on average, but cannot adapt to specific deployment conditions. In the future, online learning and adaptation techniques such as those presented in [34] can help alleviate these types of errors.
Further errors that we noticed in deployments that are not captured by our model are issues such as major thruster imbalances (housings that do not have streamlined geometry severely hinder thruster performance in asymmetric ways).
@@ V Conclusion
In this work, to our knowledge, we provide the first implementation of a highly parallelized simulation for underwater vehicle control that can be used in reinforcement learning pipelines. We use it to train, in minutes, a control policy that maps high-level 6-DOF commands directly into thruster allocations. We use domain randomization methods to address sim-to-real robustness across various dynamic settings, allowing deployments with no additional tuning at test-time. We validate this by deploying a trained policy zero-shot, running on-board in real-time a custom AUV that can recover from aggressive disturbances, with no a priori information about precise physical AUV parameters such as centers of mass and buoyancy. This approach opens the door for rapid experimentation with various other AUV configurations, without the need to design individual controllers for each.
In the future we plan to explore using online learning and adaptation techniques such as those proposed by [35, 34, 36, 37] to develop a controller that can adapt to new physical circumstances in real-time, and resolve issues such as the steady state errors we observed. Finally, we hope to integrate other types of hydrodynamics models to enable actuation methods outside of thrusters, which are substantially more difficult to develop controllers due to their highly nonlinear nature.
@@ ACKNOWLEDGMENTS
This work was supported in part by WHOI Summer Student Fellowship Program, WHOI Independent Study Award, WHOI Academic Programs Office Endowed Funds, and NSF NRI Award No. 2133029. Thanks to Ethan Fahnestock for the initial simulator and learning pipeline. Thanks also to Nathan McGuire, Peter Werner, Daniel Yang, and Tim Seyde for their help and discussions.